OpenStudy (anonymous):

Mojo jojo is standing 25.0m underneath his new mind-control machine , about to take over townsvile. but bubbles sends mojo flying straight upwards into the machine. I n order to destroy the machine, mojo must hit it w/t a velocity of atleast 30.0 m/s (u). What is the minimum initial bubbles must import to mojo.

OpenStudy (xishem):

Another kinematics question. First, write out each quantity that you know and the quantity that you are looking for. Then look for an equation that will work for those quantities.

OpenStudy (anonymous):

But I dont know how to do problems with 2 objects.

OpenStudy (anonymous):

I know that for mojo the displacment is 25 and v2 is 30.0 and thats all

OpenStudy (xishem):

The bubbles aren't really an object you need to worry about here. They just impart some initial velocity. Don't worry about them. Just know you are looking for an initial velocity. You are forgetting one quantity you know. They are on Earth, what other kinds of quantities do we know when we are in a system on Earth?

OpenStudy (anonymous):

g

OpenStudy (anonymous):

and bubbles is the name of another perosn, i think so it is a two object Q

OpenStudy (xishem):

All "Bubbles" is doing is throwing him up. Once he has given him that initial velocity, Bubbles is irrelevant to the problem. Just look at Mojo for now. These are the quantities we are dealing with:\[\Delta x=+25.0m\]\[v_f=+30.0\frac{m}{s}\]\[a_x=-9.81\frac{m}{s^2}\]\[v_i=?\]Any equations pop out? Remember that:\[\Delta x = x_f-x_i\]

OpenStudy (anonymous):

Ohh yeah... let me solve

OpenStudy (anonymous):

but when you plug it in..u get the equation equal to 5 which is wrong

OpenStudy (xishem):

Which equation are you using?

OpenStudy (anonymous):

the one u provided!